OpenStudy (anonymous):

Find the area of a regular hexagon with sides 12 ft

OpenStudy (anonymous):

A = 1/2 ap

OpenStudy (anonymous):

Any takers!

OpenStudy (anonymous):

Help me....

OpenStudy (radar):

I'm thinking lol

OpenStudy (radar):

We need to calculate height: The height is \[\sqrt{12^{2}-6^{2}}=\sqrt{144-36}=\sqrt{108}\]

OpenStudy (anonymous):

oh thank you radar.... I think this is the last one! Cause I quit! LOL

OpenStudy (radar):

Area for a regular hexagon is: A= 1/2 (6sh) A=.5 x6 x 12 x sqrt 108=374.123 sq ft

OpenStudy (radar):

s is side length, h is height A is area.

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

you gots it?

OpenStudy (anonymous):

shhhh... no lol

OpenStudy (amistre64):

hex=6 sides right?

OpenStudy (anonymous):

yeppers

OpenStudy (amistre64):

360/6 = 60 degrees for central angles then and an outer side at 12; law of cosines modified will give us lengths of sides

OpenStudy (amistre64):

12^2 = 2s^2 - 2(s)(s)(cos(60))

OpenStudy (amistre64):

144 = 2s^2 (1 - cos(60)) 144/ (1 - cos(60)) = 2s^2 144/ (1 - cos(60)) ---------------- = s^2 2 s = sqrt(all that) lol

OpenStudy (anonymous):

yikes not diggin the cos stuff....

OpenStudy (amistre64):

6.0726801....

OpenStudy (amistre64):

all that does is give me a side to go off of to caluclate the area of one tri and then gonna multiply by 6 for the whole thing

OpenStudy (anonymous):

oh danger... I think the answer is going to be ?\[\sqrt{?}\]

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

it might be :)

OpenStudy (anonymous):

is that 6.0726801 the area?

OpenStudy (amistre64):

thats the length of a side

OpenStudy (anonymous):

ewe

OpenStudy (amistre64):

we got an isoceles trianlgle with a base of 12 and a top angle of 60

OpenStudy (amistre64):

if I know the length of the sides I can caluclate the area for on tri pieces of the hex

OpenStudy (anonymous):

216 square root 3

OpenStudy (amistre64):

12^2 = s^2 + s^2 - 2ss cos(60) 144 = 2s^2 - 2s^2 cos(60) 144 = 2s^2(1 - cos(60)) 144 ---------- = 2s^2 (1-2cos(60)) 72 ---------- = s^2 (1-2cos(60)) 72 sqrt[ ----------- ] = s (1-2cos(60))

OpenStudy (amistre64):

s*s*sin(60) ---------- = area of tri piece; 6 of those is the whole thing right? 2

OpenStudy (anonymous):

right cause its a 6 sided polygon?!

OpenStudy (amistre64):

sqrt*sqrt = innards

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

that 2cos is a typo.... that denom is sposed to be 1-cos(60)

OpenStudy (amistre64):

72 sin(60) ------------ = area of tri 2 (1-cos(60)) 36 sin(60) ----------- (1-cos(60)) 36 (sqrt(3)/2) ------------- (1-(1/2))

OpenStudy (amistre64):

18 sqrt(3) --------- = 36sqrt(3) right? 1/2

OpenStudy (amistre64):

times 6 = 216 sqrt(3)

OpenStudy (anonymous):

lol here... A = 1/2 ap A= 1/2 (6sqrt3)(72) 216sqrt3 = 374.1cm sq

OpenStudy (anonymous):

YOU GOT IT!!!!!!!!!!!!!

OpenStudy (amistre64):

I did lol

OpenStudy (anonymous):

and it matches mine :) Good Business! You soooo hired!! LOL

OpenStudy (anonymous):

are

OpenStudy (amistre64):

yay!!.... and i did it all with one brain cell tied behind my cerebellum ;)

OpenStudy (anonymous):

Hilarious!! You rock!

OpenStudy (amistre64):

gotta go write some homework.... have fun ;)

OpenStudy (anonymous):

thank you!

OpenStudy (radar):

I was right, he got the same answer lol

OpenStudy (radar):

How did it become centimeters, I thought it was 12 ft

OpenStudy (anonymous):

ft squared lol

OpenStudy (radar):

Great, good luck with your studies.

OpenStudy (anonymous):

ha thanks